OpenStudy (anonymous):

pls help step by step 1/2x = 10

OpenStudy (saifoo.khan):

x = 20

OpenStudy (saifoo.khan):

simply cross multiply.

OpenStudy (llort):

Possible intermediate steps: 1/(2 x) = 10 Divide both sides by 1/2: 1/x = 20 Raise both sides to the power of -1: x = 1/20

OpenStudy (anonymous):

Multiply: ?

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (llort):

saifoo... I think it is actually 1/20

OpenStudy (saifoo.khan):

umm, nope.

OpenStudy (llort):

you multiply by 2 and you are left with 1/x=20

OpenStudy (llort):

I still disagree...

OpenStudy (saifoo.khan):

wait, let me rewrite the Prob.

OpenStudy (llort):

if it is 1/2x, then the x is in the denominator, not the numerator.

OpenStudy (anonymous):

it says Multiply: b. 1/4s = 8

OpenStudy (saifoo.khan):

is the prob like as shown in the Picture?

OpenStudy (anonymous):

yes

OpenStudy (llort):

so is it 1/4(s) or 1/4s?

OpenStudy (saifoo.khan):

then im Correct.

OpenStudy (llort):

because it was written wrong...

OpenStudy (anonymous):

its 1/4s